Suppose you have a near-point distance of 26 cm. The angular resolution of the eye is determined by diffraction (see the Rayleigh criterion) and is approximately 1 arcmin. (Note that 1°=60 arcmin.) What is the smallest object your bare eye can resolve?
Answer #1

1 arcmin =0.290888* 10^-3 rad

the minimum separation between object is

theta_min = d/L

d = theta_min * L

= 0.290888* 10^-3 rad ( 26 cm) = 7.56 * 10^-3 cm